OpenStudy (anonymous):

A father is 1 year older than 3 times the age of his son. The son is 20 years old. Let F = the age of the father. Which formula would calculate the father's age?

OpenStudy (anonymous):

more options please... unless your supposed to write it yourself??

OpenStudy (anonymous):

and by it i mean the equation

OpenStudy (mathstudent55):

Let s = age of son F = 3s + 1 s = 20 F = 3(20) + 1

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

s=20 f-1=3s f=3*20+1=61

OpenStudy (anonymous):

1. F=3(20-1) 2. F + 1 = 3(20) 3. F - 1 = 3(20) these are the three i have to pick from

OpenStudy (mathstudent55):

Look at the last one: F - 1 = 3(20) add 1 to both sides: F = 3(20) + 1 That was my solution